OpenStudy (anonymous):

A gene can be either type A or type B, and it can be either dominant or recessive. If the gene is type B, then there is a probability of 0.31 that it is dominant. There is also a probability of 0.22 that a gene is type B and it is dominant. What is the probability that a gene is of type A? .......This question is killing me.

OpenStudy (zarkon):

given: \[P(D|B)=.31\] \[P(B,D)=.22\] so \[.31=P(D|B)=\frac{P(B,D)}{P(B)}=\frac{.22}{P(B)}\] thus \[P(B)=\frac{.22}{.31}\approx.709677419355\] \[P(A)=1-P(B)\approx1-709677419355=.290322580645\]